# Neural Network ANN Artificial Neural Network

### v1

@@ Abstract
Table tennis is a fast-paced and exhilarating sport that demands agility, precision, and fast reflexes. In recent years, robotic table tennis has become a popular research challenge for robot perception algorithms. Fast and accurate ball detection is crucial for enabling a robotic arm to rally the ball back successfully. Previous approaches have employed conventional frame-based cameras with Convolutional Neural Networks or traditional computer vision methods. In this paper, we propose a novel solution that combines an event-based camera with Spiking Neural Networks for ball detection. We use multiple state-of-the-art SNN frameworks and develop a SNN architecture for each of them, complying with their corresponding constraints. Additionally, we implement the SNN solution across multiple neuromorphic edge devices, conducting comparisons of their accuracies and run-times. This furnishes robotics researchers with a benchmark illustrating the capabilities achievable with each SNN framework and a corresponding neuromorphic edge device. Next to this comparison of SNN solutions for robots, we also show that an SNN on a neuromorphic edge device is able to run in real-time in a closed loop robotic system, a table tennis robot in our use case.
@@ I INTRODUCTION
In recent years, robotic table tennis has become a popular research challenge for perception and robot control. While not yet able to compete with professional players, table tennis robots are an exciting research environment to bring perception and control algorithms towards their limits [1]. Fast and accurate ball detection is a crucial perception task for a table tennis robot system. So far, most research uses frame-based cameras together with a Convolutional Neural Network (CNN) based ball detection or a classical computer vision approach [2][3][4][5]. In this work, we enter a different avenue and combine two more recent technologies, event-based cameras and Spiking Neural Networks (SNNs).
@@
SNNs mimic the spiking behavior of biological neurons, offering a biologically inspired approach to Neural Network (NN) computation. Unlike traditional artificial neurons that produce real-valued outputs, spiking neurons receive input spikes and integrate them in the state of the neuron, called membrane potential.
@@
Figure 1: (left) Five observed 2D trajectories in the camera frame of the event-based camera with ground truth in green and the estimated positions in red. (right) An example of a single 2D ball detection in a region of interest of $64$x64646464 pixel of an accumulated event frame which serves as the input to the Spiking Neural Network (SNN)) with ground truth in green and the estimated position in red.
When the neuron’s membrane potential reaches a defined threshold, the neuron emits a spike that propagates through the network, and resets its membrane potential. This binary output format of SNNs enables energy-efficient computations on neuromorphic hardware, since no multiplications are needed [6]. Neuromorphic hardware, specifically designed for efficient SNN processing, can leverage the sparsity and binary nature of SNN outputs.
@@
However, efficient processing relies on gathering suitable data. Conventional cameras capture frames at a fixed frame rate, providing information about brightness and color. Event-based cameras, on the other hand, report asynchronous brightness changes per pixel, without measuring the absolute brightness [7] [8]. These cameras offer a high dynamic range, a temporal resolution in the order of $\mu$s, energy and data efficiency. The binary output of event-based cameras aligns with the spike format of SNNs, making them a good match.
@@
While ball detection solutions with frame-based cameras are successfully used, the high temporal resolution of event-based cameras promises faster and more frequent ball detections. This can improve the prediction of the ball’s trajectory, which allows a faster robot control. SNNs align well with event data and can handle more complex scenarios, like a cluttered environment, compared to model-based solutions.
@@
This work explores the combination of an event-based camera and SNNs for object detection and their deployment on neuromorphic edge devices. We report the error and run-time, both in simulation and on multiple neuromorphic edge devices, namely the DynapCNN [^1] https://www.synsense.ai/products/dynap-cnn/ from SynSense, Akida [^2] https://brainchip.com/akida-enablement-platforms/ from BrainChip and Loihi2333https://www.intel.com/content/www/us/en/research/neuromorphic-computing-loihi-2-technology-brief.html from Intel. Benchmarking these three widely used neuromorphic edge devices on our use case of object detection gives an example of what is possible with these edge devices in robotics perception applications.
Contributions of this work are as follows:
  • We show how event-based cameras and SNNs complement each other
  • We present an object detection system for fast-moving objects, capable of running in real-time
  • We show the capabilities and limitations of a deployment on three neuromorphic edge devices
  • We deploy an SNN for object detection on a neuromorphic edge device in a real-time table tennis robot system
  • We make the used dataset publicly available
@@ II-B Spiking Neural Networks (SNN)
Studies indicate that SNNs are more energy-efficient [14] and excel in processing spatio-temporal data due to their use of sparse, binary signals [17]. This efficiency stems from requiring only simple additions rather than complex multiplications, reducing both computational and memory demands [6]. Consequently, SNNs are ideal for energy-sensitive hardware applications [18][19].
A significant advancement in SNN research is the development of efficient training techniques. Converting pre-trained Artificial Neural Networks into SNNs, will be described in Section II-C and direct training of SNNs using spike-based learning rules, in Section II-D.
@@ II-C Artificial Neural Network (ANN) to SNN Conversion
One approach to train SNNs involves converting pre-trained ANNs into SNNs [20].
This can be done by approximating the output of the Rectified Linear Unit (ReLU) non-linearity with rate-code [21]. Rate-code represents the relative frequency of spikes, obtained by replacing the ReLU with the Heaviside step function and setting a spiking threshold of one. This approximation captures the essential characteristics of the ReLU, for output activities between zero and one. ReLU outputs larger than one can not be represented by normal IF neurons, as they can at most spike once per time step. However, a multi-spike neuron model [22] allows the direct representation of the entire ReLU output space through rate-code.
@@ II-E Spiking Object Detection
Researchers have explored using SNNs for specific object detection tasks. In [28], SNNs were employed to predict ball trajectories using data from event-based cameras. Their approach utilized spatio-temporal filters with weights incorporating delays to capture temporal information. They also employed leaky IF neurons and trained the network using STDP.
[29] utilizes an SNN for detecting balls and pipes in a simulated environment. They used simulated neuromorphic data and designed an SNN architecture specifically tailored for detecting circles, ellipses, and lines using the Hough transform. Their approach demonstrated successful detection using the designed SNN without training, relying on hard-coded algorithms.
@@ III METHOD
We recorded a custom-made dataset to train our SNNs, which we describe in Section III-A. In this work, we used three state-of-the-art SNN frameworks and a corresponding neuromorphic edge device. For each of them, we designed an SNN architecture, conforming with their specific constraints, described in Section III-B. The training with the three SNN frameworks is covered in Section III-C.
@@ III-A1 Recording setup
To calibrate this camera system containing frame- and event-based cameras with baselines between $3$m and $5$m, we used the wand-based calibration approach introduced in [30]. The frame-based cameras are used to detect and triangulate the position of the table tennis ball in 3D, as described in [2]. The event stream from the event-based cameras is used as input for our SNN approach.
A Butterfly Amicus Prime ball gun, with default speed settings ($4$m/s), was used to shoot balls in front of a static background.
@@ III-A2 Data generation
The input of our SNN is a matrix of size $64$x64646464 pixels, where over a time range of $1$ms the pixel value is set to one if at least one event occurred at the pixel and zero otherwise. This event representation removes information about the event’s polarity. Keeping both polarities would require multiple input channels and, therefore, a bigger network. To achieve fast inference times, we favored smaller networks. As in [31], this event representation does also not differentiate between pixels with more or less events, as long as they have any. This reduction in information was carried out as it led to better performance in our experiments. A dynamic Region of Interest (ROI) was used to crop $64$x64646464 pixels out of the $1280$x720720720720 pixels output of the event-based camera, similar to [2]. This was done by setting the last known ball position as the center of the ROI. This has the benefit of being applicable in real-time using the system’s own past ball position estimates. Since the system starts each trajectory without past position estimates, we used manually defined start regions. This limitation could be alleviated by using the position of a frame-based detection system or an CNN based ball detector for the initialization.
The output of the SNN is the 2D pixel position of the detected ball. Therefore, the required ground truth is the 2D ball position in the image frame of the event-based camera. We used two ways to generate the 2D ball positions, serving as ground truth. We projected the triangulated 3D ball positions, which we get from the frame-based cameras using [2] into the camera frame of the event-based camera. This way, we generated $7569$ ground truth positions. Since we only obtain 3D positions every $7$ms, we additionally labeled the 2D ball position manually on an additional set of $2123$ event frames. An example is shown in Fig. 3.
Figure 3: An example of a manually labeled 2D ball position on an accumulated event frame ($1$ms accumulation time). Accumulated event frames are generated by indicating if an event occurred per pixel over a defined time interval.
@@ III-B Network Architecture
We cast the task of detecting the table tennis ball in 2D as a classification task. This allows fewer time steps and, therefore, a faster inference time than a regression network due to the rate-code approximation. We treat the $x$- and $y$-positions as two independent classification tasks. In this case, the ball’s $x$-position can be one of $64$ classes, as can the $y$-position. A visualization of the network’s output is shown in Fig. 4. Although the neuromorphic edge devices support higher input resolutions ($128$x128128128128 pixels for the DynapCNN), we decided to use only $64$x64646464 pixels to improve the inference time, which is crucial for our real-time use case. To receive the network’s prediction, the neurons with the maximal output are determined in each of the two populations.
@@
TABLE II: Network architecture for the three SNN frameworks. The activation function and the pooling are given in the following order: sinabs / MetaTF / Lava. “None” indicates that no layer is used for the corresponding SNN framework. Where stride is not explicit, it defaults to $1$.
@@ III-B2 MetaTF (Akida)
The network we used to train with the MetaTF framework uses batch normalization layers in all layers except the last one, as it led to improved performance, and the Brainchip supports the use of biases, which are necessary to integrate batch normalization into an SNN without significant effect on the inference time. Quantized ReLUs are used because the Akida chip squashes the rate-code approximation of the ReLU into one time step, where it is then represented by a step-wise quantized ReLU. In the first layer, we use max pooling instead of average pooling, since this is the pooling type supported by the MetaTF framework. For the use on the chip, pooling comes with additional restrictions, which is why there is no second pooling layer and instead a stride of two in the second convolutional layer.
@@ III-C Training
For all SNN frameworks, we trained the networks on the GPU and deployed them on the edge devices for inference. We used all the $7569$ automatically generated 2D ground truth positions and additional $1061$ manually labeled positions (as described in Section III-A2) for the training set and $531$ manually labeled 2D positions for the validation and the test set.
Since the training depends on the SNN framework, we describe the training for each one. Common hyperparameters are listed in Table III.
@@ III-C1 Training with sinabs (DynapCNN)
To minimize the quantization loss, the distribution of weights needs to be such that no extreme outlier weights exist which would decrease the quantization accuracy for all other weights. To normalize the weights on the DynapCNN, we used an additional loss term using only the absolute maximum of the weights for every layer.
The SNN was trained using the periodic exponential function as surrogate function [25], with back-propagation through time (BPTT).
@@ IV EXPERIMENTS
We performed two sets of experiments. In the first one, described in Section IV-A, we evaluate the detection error and measure the time per forward pass of the three different SNN frameworks in simulation as well as on the neuromorphic edge device. Section IV-B describes the second experiment, in which we run the SNN on a neuromorphic edge device within a table tennis robot setup in real-time.
Figure 5: (left) Experiment setup for the offline experiment. (right) Experiment setup for the online experiment.
@@ IV-A Offline experiment
In this experiment, we compare the different SNN frameworks in terms of error and time per forward pass given our robotic perception task. We recorded data with the setup introduced in Section III-A. Given this recorded data, we ran the 2D ball detection $10$ times for BrainChip and sinabs and $5$ times for Lava. The setup is depicted in Fig. 5 on the left. Five recorded 2D ball trajectories and a single 2D ball detection are shown in Fig. 1.
@@
Steps indicate how many time steps were used for the rate-code for one inference pass, a trade-off between speed and accuracy. The steps count in brackets represents the number of time steps when using multispike neurons.
As can be seen in Table IV the BrainChip framework has the lowest ball center error of $1.44$ pixels. The error of sinabs and Lava are $1.55$ pixels and $1.57$ pixels, respectively, quite close together. These errors are comparable with the frame-based ball detection approach in [34], which achieves an error of $1.32$ pixels.
Important to note is that depending on the neuromorphic edge device, the total time per forward pass is heavily influenced by the integration into the system. The total time for a forward pass includes the inference time and, depending on the neuromorphic edge device data transfer, pre- and post-processing. Therefore, we list the total time per forward pass and the inference time in Table V.
TABLE V: Total time per forward pass and inference time of the different neuromorphic edge devices with mean and std. dev over $10$ runs. $*{}^{*}$rocessing one input on the DynapCNN takes an average of $1.765$ms, but with delays from the USB connection, the time for a forward pass is significantly higher.
Edge device One forward pass [ms] Inference time [ms]
BrainChip Akida 2.20 $\pm$ 0.35 0.89 $\pm$ 0.28
DynapCNN 46.04 $\pm$ $21.38^{*}$ 0.82 $\pm$ 0.24
Loihi2 1458.57 $\pm$ 230.50 0.62 $\pm$ 0.01
@@
For Loihi2, we did not have direct access to an edge device, but only through a virtual machine provided by Intel via the Intel Neuromorphic Research Community [^8] https://intel-ncl.atlassian.net/wiki/spaces/INRC/overview. With this Loihi2 setup, the time per forward pass is with $1458.57$ms two orders of magnitude slower than the other devices. This is mainly due to inefficient communication between the virtual machine and the Loihi2 chip, which is a consequence of our current implementation not being well optimized in this regard.
Overall, we can see that while the errors of the different neuromorphic frameworks are quite close together, the difference in inference time and the time per forward pass varies significantly. As already mentioned, the way in which the neuromorphic edge devices are connected to the PC plays a crucial role.

### v2

# Neuromorphic Computing NN Neural Network

@@ Abstract
Neuromorphic Computing (NC) and Spiking Neural Networks in particular are often viewed as the next generation of Neural Networks. NC is a novel bio-inspired paradigm for energy efficient neural computation, often relying on SNNs in which neurons communicate via spikes in a sparse, event-based manner. This communication via spikes can be exploited by neuromorphic hardware implementations very effectively and results in a drastic reductions of power consumption and latency in contrast to regular GPU-based NNs. In recent years, neuromorphic hardware has become more accessible, and the support of learning frameworks has improved. However, available hardware is partially still experimental, and it is not transparent what these solutions are effectively capable of, how they integrate into real-world robotics applications, and how they realistically benefit energy efficiency and latency. In this work, we provide the robotics research community with an overview of what is possible with SNNs on neuromorphic hardware focusing on real-time processing. We introduce a benchmark of three popular neuromorphic hardware devices for the task of event-based object detection. Moreover, we show that an SNN on a neuromorphic hardware is able to run in a challenging table tennis robot setup in real-time.
@@ I INTRODUCTION
Spiking Neural Networks (SNNs) mimic the spiking behavior of biological neurons, offering a biologically inspired approach to Neural Network (NN) computation. Unlike traditional artificial neurons that produce real-valued outputs, spiking neurons receive input spikes and integrate them in the state of the neuron, called membrane potential.
@@
Figure 1: Left: Three examples of 2D ball detections in an accumulated event frame which serves as the input to the Spiking Neural Network (SNN) with ground truth in green and the estimated position in red. Right: Five observed 2D trajectories in the camera frame of the event-based camera with ground truth in green and the estimated positions in red. Background: The table tennis robot setup with the robot hitting back a table tennis ball in a rally.
When the neuron’s membrane potential reaches a defined threshold, the neuron emits a spike that propagates through the network, and resets its membrane potential. More theoretical work on SNNs has already been present in research for decades [1][2][3][4]. With the availability of more computing power through GPUs, it became feasible to train and simulate SNNs for real-world tasks. However, simulating SNNs is very inefficient on GPUs since the Neuromorphic Computing (NC) paradigm is fundamentally different. Therefore, running SNNs on GPUs, though possible, is not a reasonable option for real-world applications, especially not for real-time robotics applications.
In contrast, neuromorphic hardware, specifically designed for efficient SNN processing, can leverage the sparsity and binary nature of SNN outputs, resulting in a drastic reduction of power consumption and latency compared to conventional NNs.
@@
But the efficient processing of SNNs relies on gathering suitable data. Conventional cameras capture frames at a fixed frame rate, providing information about brightness and color. Event-based cameras, on the other hand, report asynchronous brightness changes per pixel, without measuring the absolute brightness [5][6]. These cameras offer a high dynamic range, a temporal resolution in the order of $\mu$s, as well as energy and data efficiency. The binary output of event-based cameras aligns with the spike format of SNNs, making them a good match.
This synergy between event-based cameras and SNNs presents an opportunity to improve real-time performance in robotics. One such area where these advancements can be applied is table tennis robotics, which has gained popularity in recent years [7]. While not yet able to compete with professional players, table tennis robots are an exciting research environment to bring algorithms towards their limits. We thus use a table tennis robot scenario as a benchmark suite for event-based neuromorphic perception systems.
@@
A primary perception task for a table tennis robot system is fast and accurate ball detection. So far, most research uses frame-based cameras together with a Convolutional Neural Network (CNN) based ball detection or a classical computer vision approach [8][9][10][11]. While ball detection solutions with frame-based cameras are successfully used, the high temporal resolution of event-based cameras promises faster and more frequent ball detections. This can improve the prediction of the ball’s trajectory, which allows a faster robot control. As mentioned, SNNs align well with event data and can handle more complex scenarios, like a cluttered environment, compared to model-based solutions.
@@
This work explores the combination of an event-based camera and SNNs for object detection and their deployment on neuromorphic edge devices. We analyze the potential benefits of this fusion, and discuss the limitations of a deployment on neuromorphic edge devices. We report the error and run-time, both in simulation and on multiple neuromorphic edge devices, namely the DynapCNN [^1] https://www.synsense.ai/products/dynap-cnn/ from SynSense, Akida [^2] https://brainchip.com/akida-enablement-platforms/ from BrainChip and Loihi2333https://www.intel.com/content/www/us/en/research/neuromorphic-computing-loihi-2-technology-brief.html from Intel. Benchmarking these three widely used neuromorphic edge devices on our use case of object detection gives an example of what is possible with these edge devices in robotics perception applications.
In summary, our contributions with this work are:
  • We demonstrate the effectiveness of SNNs for event-based object detection
  • We conduct a comparative study of available neuromorphic edge devices
  • We present a streamlined integration of event-based sensing, neuromorphic processing, and robot arm planning and control in a table tennis robot setup
  • We provide a publicly available benchmark dataset for event-based ball detection
@@ II-B Spiking Neural Networks (SNN)
Studies indicate that SNNs are more energy-efficient [17] and excel in processing spatio-temporal data due to their use of sparse, binary signals [20]. Consequently, SNNs are ideal for energy-sensitive hardware applications [21][22].
A significant advancement in SNN research is the development of efficient training techniques. Converting pre-trained Artificial Neural Networks into SNNs will be described in Section II-C and direct training of SNNs using spike-based learning rules, in Section II-D.
@@ II-C Artificial Neural Network (ANN) to SNN Conversion
One approach to train SNNs involves converting pre-trained ANNs into SNNs [23]. This can be done by approximating the output of the Rectified Linear Unit (ReLU) non-linearity with rate-code [24]. Rate-code represents the relative frequency of spikes, obtained by replacing the ReLU with the Heaviside step function and setting a spiking threshold of one. This approximation captures the essential characteristics of the ReLU, for output activities between zero and one. ReLU outputs larger than one can not be represented by normal IF neurons, as they can at most spike once per time step. However, a multi-spike neuron model [25] allows the direct representation of the entire ReLU output space through rate-code.
@@ II-E Spiking Object Detection
SNNs were also explored for specific object detection tasks. In [31], SNNs were employed to predict ball trajectories using data from event-based cameras. Their approach utilized spatio-temporal filters with weights incorporating delays to capture temporal information. They also employed leaky IF neurons and trained the network using STDP.
An SNN was utilized in [32] for detecting balls and pipes in a simulated environment. They used simulated neuromorphic data and designed an SNN architecture specifically tailored for detecting circles, ellipses, and lines using the Hough transform. Their approach demonstrated successful detection using the designed SNN without training, relying on hard-coded algorithms.
@@ III METHOD
We recorded a ball detection dataset to train our SNNs, described in Section III-A. In this work, we used three state-of-the-art SNN frameworks and a corresponding neuromorphic edge device. For each of them, we designed an SNN architecture, conforming with their specific constraints, described in Section III-B. The training with the three SNN frameworks is covered in Section III-C.
@@ III-A1 Recording setup
To calibrate this camera system containing frame- and event-based cameras, we used the wand-based calibration approach introduced in [33]. The frame-based cameras are used to detect and triangulate the position of the table tennis ball in 3D, as described in [8]. The event stream from the event-based cameras is used as input for our SNN approach.
A Butterfly Amicus Prime ball gun, with default speed settings ($4$m/s), was used to shoot balls, as shown in the background of Fig. 1.
@@ III-A2 Data generation
The input of our SNNs is a matrix of size $64$x64646464 pixels, where over a time range of $1$ms the pixel value is set to one if at least one event occurred at the pixel and zero otherwise. To reduce computation and achieve faster inference times, we favored smaller networks and thus did not include the polarity information of the events. As in [34], our event representation does also not differentiate between pixels with more or less events, as long as they have any. This reduction in information was carried out as it led to better performance in our experiments. A dynamic Region of Interest (ROI) was used to crop $64$x64646464 pixels out of the $1280$x720720720720 pixels output of the event-based camera, similar to [8]. This was done by setting the last known ball position as the center of the ROI. This has the benefit of being applicable in real-time using the system’s own past ball position estimates. Since the system starts each trajectory without past position estimates, we used manually defined start regions. This limitation could be alleviated by using the position of a frame-based detection system or an CNN based ball detector for the initialization.
The output of the SNNs is the 2D pixel position of the detected ball. Therefore, the required ground truth is the 2D ball position in the image frame of the event-based camera. We used two ways to generate the 2D ball positions, serving as ground truth. We projected the 3D ball positions from the frame-based camera system into the camera frame of the event-based cameras. This way, we generated $7569$ ground truth positions. Since we only obtain 3D positions every $7$ms, we additionally labeled the 2D ball position manually on an additional set of $2123$ event frames.
@@ III-B Network Architecture
We cast the task of detecting the table tennis ball in 2D as a classification task. This allows fewer time steps and, therefore, a faster inference time than a regression network due to the rate-code approximation. We treat the $x$- and $y$-positions as two independent classification tasks. In this case, the ball’s $x$-position can be one of $64$ classes, as can the $y$-position. A visualization of the network’s output is shown in Fig. 3. Although the neuromorphic edge devices support higher input resolutions, we decided to use only $64$x64646464 pixels to improve the inference time, which is crucial for our real-time use case. To receive the network’s prediction, the neurons with the maximal output are determined in each of the two populations.
@@
TABLE II: Network architecture for the three SNN frameworks. The details are given in the following order: sinabs / MetaTF / Lava. “None” indicates that no layer is used for the corresponding SNN framework. Where stride is not explicit, it defaults to $1$.
@@ III-B2 MetaTF (Akida)
The network we used to train with the MetaTF framework uses batch normalization layers in all layers except the last one, as it led to improved performance, and Akida supports the use of biases, which are necessary to integrate batch normalization into an SNN without significant effect on the inference time. Quantized ReLUs are used because the Akida chip squashes the rate-code approximation of the ReLU into one time step, where it is then represented by a step-wise quantized ReLU. In the first layer, we use max pooling instead of average pooling, since this is the pooling type supported by the MetaTF framework. For the use on the chip, pooling comes with additional restrictions, which is why there is no second pooling layer and instead a stride of two in the second convolutional layer.
@@ III-C Training
For all SNN frameworks, we trained the networks on a GPU and deployed them on the edge devices for inference. We used all the $7569$ automatically generated 2D ground truth positions and additional $1061$ manually labeled positions (as described in Section III-A2) for the training set and $531$ manually labeled 2D positions for the validation and the test set. Since the training depends on the SNN framework, we describe the training for each one. Common hyperparameters are listed in Table III.
@@ III-C1 Training with sinabs (DynapCNN)
To minimize the quantization loss, the distribution of weights needs to be such that no extreme outlier weights exist, which would decrease the quantization accuracy for all other weights. To normalize the weights on the DynapCNN, we used an additional loss term using only the absolute maximum of the weights for every layer. The SNN was trained using the periodic exponential function as surrogate gradient function [28] and backprop through time.
@@ IV EXPERIMENTS
In the first experiment, described in Section IV-A, we evaluate the detection error and measure the time per forward pass of the three different SNN frameworks in simulation as well as on the neuromorphic edge device. Section IV-B describes the second experiment, in which we run the SNN on a neuromorphic edge device within a table tennis robot setup in real-time.
Figure 4: Left: Experiment setup for the offline experiment. Right: Experiment setup for the online experiment.
@@ IV-A Offline experiment
In this experiment, we compare the different SNN frameworks in terms of error and time per forward pass, given our robotic perception task. Using our recorded data, we ran the 2D ball detection $10$ times for BrainChip and sinabs and $5$ times for Lava. The setup is depicted in Fig. 4 on the left. Five recorded 2D ball trajectories and three 2D ball detections are shown in Fig. 1.
@@
TensorFlow RTX 2080 Ti - 3.55 $\pm$ 1.49 3.49 $\pm$ 0.62
Steps indicate how many time steps were used for the rate-code for one inference pass, a trade-off between speed and accuracy. The steps count in brackets represents the number of time steps when using multispike neurons. We also report the results of an existing setup [37], based on [10], as a comparison. This setup uses a frame-based camera and a CNN with similar image resolution and number of convolutional layers on an RTX 2080 Ti GPU.
As can be seen in Table IV the BrainChip framework has the lowest ball center error of $1.44$ pixels. The error of sinabs and Lava are $1.55$ pixels and $1.57$ pixels, respectively, quite close together.
Important to note is that depending on the neuromorphic edge device, the total time per forward pass is heavily influenced by the integration into the system. The total time for a forward pass includes the inference time and, depending on the neuromorphic edge device data transfer, pre- and post-processing. Therefore, we list the total time per forward pass, the inference time, and the power consumption in Table V.
TABLE V: Total time per forward pass and inference time of the different neuromorphic edge devices with mean and std. dev over $10$ runs and the power consumption. Processing one input on the DynapCNN takes an average of $1.765$ms, but with delays from the USB connection, the time for a forward pass is significantly longer. Taken from specs.
Device One forward pass [ms] Inference time [ms] Power consumption [mW]
Akida 2.20 $\pm$ 0.35 0.89 $\pm$ 0.28 $\sim 4.5$
DynapCNN 46.04 $\pm$ 1.38 0.82 $\pm$ 0.24 $\sim 5^{*}$
Loihi2 1458.57 $\pm$ 230.50 0.62 $\pm$ 0.01 $\sim 100^{*}$
RTX 2080 Ti 3.49 $\pm$ 0.62 1.77 $\pm$ 0.28 $\sim 250000^{*}$
@@
For Loihi2, we did not have direct access to an edge device, but only through a virtual machine provided by Intel via the Intel Neuromorphic Research Community [^8] https://intel-ncl.atlassian.net/wiki/spaces/INRC/overview. With this Loihi2 setup, the time per forward pass is with $1458.57$ms two orders of magnitude slower than the other devices. This is mainly due to inefficient communication between the virtual machine and the Loihi2 chip, which is a consequence of our current implementation not being well optimized in this regard. Note that the inference time on the GPU is longer than on all neuromorphic edge devices, however, the good hardware integration of the GPU shows its advantage.
Regarding the power consumption, we measured the one for Akida, but had to rely on specs for the other devices. The power consumption for Akida and DynapCNN is quite close together. Loihi2, being a much more powerful and versatile platform, consumes an order of magnitude more energy, still several orders of magnitude lower than a GPU, showing the benefits of Neuromorphic Computing (NC) on neuromorphic edge devices for robotic applications.
Overall, we can see that while the errors of the different neuromorphic frameworks are similar, the difference in inference time and the time per forward pass varies significantly. As already mentioned, the way in which the neuromorphic edge devices are connected to the PC plays a crucial role.